# Courcelle’s Theorem for Lipschitz Continuity

### v1

@@ 2.3 Monadic Second-Order Logic
Then, we can design an algorithm for an (C)MSO maximization (minimization) problem over graphs of bounded treewidth. For simplicity, we assume that the given graph $G$ has no sources, that is, $\mathsf{src}(G)=\varnothing$. Let $t$ be a term denoting $G_{t}$ and $\varphi(X)$ be the (C)MSO-formula describing the constraint of the problem. We recursively compute the value $\mathsf{opt}(t^{\prime},\varphi^{\prime}\mathord{{\restriction}_{S}})=\max\{w(A):G\models\varphi^{\prime}\mathord{{\restriction}_{S}}(A)\}$ for any subterm $t^{\prime}$ of $t$ and MSO-formula $\varphi^{\prime}\mathord{{\restriction}_{S}}$ as follows, where $w\in\mathbb{R}_{\geq 0}^{V}$ is the given weight vector and $w(X)=\sum_{x\in X}w_{x}$ for any $X\subseteq V(G)$. If $t^{\prime}$ is of the form $\mathsf{fg}_{B}(t^{\prime\prime})$, then $\mathsf{opt}(\mathsf{fg}_{B}(t^{\prime\prime}),\varphi^{\prime}\mathord{{\restriction}_{S}})=\max_{S^{\prime}\subseteq B}\{\mathsf{opt}(t^{\prime\prime},\psi\mathord{{\restriction}_{S\cup S^{\prime}}})+w(S^{\prime})\}$, where, $\psi$ is the formula obtained from Corollary 2.6. If $t^{\prime}$ is of the form $t_{1}\mathbin{//}t_{2}$, then $\mathsf{opt}(t^{\prime},\varphi^{\prime}\mathord{{\restriction}_{S}})=\max_{i\in[p]}\{\mathsf{opt}(t_{1},\theta\mathord{{\restriction}_{S_{i}}})+\mathsf{opt}(t_{2},\psi\mathord{{\restriction}_{S^{\prime}_{i}}})\}$, where, $\{\theta_{i}\mathord{{\restriction}_{S_{i}}}({X}),\psi_{i}\mathord{{\restriction}_{S^{\prime}_{i}}}({X})\}_{i\in[p]}$ are the formulas obtained from Corollary 2.6. Then, $\mathsf{opt}(t,\varphi\mathord{{\restriction}_{\varnothing}})$ is the maximum value for the problem since $t$ has no sources. In the following, for simplicity, we may omit the $\mathord{{\restriction}_{S}}$ notation and consider only formulas $\varphi$ such that $\mathsf{sat}(G,\varphi)$ does not contain any sources.
In the above setting, the graphs are considered without any special vertex sets (called as colors) or special vertices (called as labels). However, for $k$-graphs with a constant number of colors and/or labels, an appropriate HR-algebra and logics can be defined, and the similar results for Corollary 2.6 hold (see e.g. [CourcelleE2012, Libkin04]).
@@ 2.4 Lipschitz Continuity
In our algorithm, we use the following procedures softmax and softmin. These procedures are derived from the exponential mechanism in the literature of differential privacy [Dwork06] and frequently appear in the literature on Lipschitz continuity [KumabeY23, KumabeY25]. Here, we organize them into a more convenient form for our use. Let $p\in\mathbb{Z}_{\geq 1}$, $x_{1},\dots,x_{p}\in\mathbb{R}_{\geq 0}$, and $\epsilon\in(0,1]$. The softmax of $x_{1},\dots,x_{p}$ is taken as follows. If $\max_{i\in[p]}x_{i}=0$, we set $\mathsf{softargmax}^{\epsilon}_{i\in[p]}x_{i}$ to be an arbitrary index $i$ with $x_{i}=0$ and $\mathsf{softmax}^{\epsilon}_{i\in[p]}x_{i}=0$. Assume otherwise. First, we sample $\bm{c}$ uniformly from $\left[\frac{2\epsilon^{-1}\log(2p\epsilon^{-1})}{\max_{i\in[p]}x_{i}},\frac{4\epsilon^{-1}\log(2p\epsilon^{-1})}{\max_{i\in[p]}x_{i}}\right]$. Let $\bm{i}^{*}$ be a probability distribution over $[p]$ such that
@@
holds for all $i\in[p]$. Then, we define $\mathsf{softargmax}^{\epsilon}_{i\in[p]}x_{i}:=\bm{i}^{*}$ and $\mathsf{softmax}^{\epsilon}_{i\in[p]}x_{i}:=x_{\bm{i}^{*}}$. We have the following. The proofs of Lemmas 2.8 and 2.9 are given in Appendix B.
Lemma 2.8.
@@
holds for all $i\in[p]$. Then, we define $\mathsf{softargmin}^{\epsilon}_{i\in[p]}x_{i}:=\bm{i}^{*}$ and $\mathsf{softmin}^{\epsilon}_{i\in[p]}x_{i}:=x_{\bm{i}^{*}}$. We have the following. The proofs of Lemmas 2.10 and 2.11 are given in Appendix B.
Lemma 2.10.
@@ 3 Lipschitz Continuous Algorithms for $\textsf{MSO}_{2}$ Optimization Problems
In this section, we prove Theorems 1.1, 1.2 and 1.3. Since the proofs are similar to that for Theorem 1.1, we provide most of the proof of Theorem 1.2 in Appendix C. In Section 3.1, we define the notations and give a brief overview of the algorithm. Section 3.2 handles the base case, where the graph consists of a single vertex. Sections 3.3 and 3.4 analyze the impact of the parallel composition and the forget operations, respectively, on approximability and the Lipschitz constant. In Section 3.5, we put everything together to prove Theorem 1.1. Finally, in Section 4.1, we provide a more refined analysis for the special case of the maximum weight independent set problem to prove Theorem 1.3.
@@ 3.1 Definitions
for minimization problems. If $\mathsf{sat}(G,\varphi)=\emptyset$, we define $\mathsf{opt}_{w}[G,\varphi]=\mathtt{nil}$. For graphs $G$ and $\textsf{MSO}_{2}$ formulas $\varphi$ with $\mathsf{sat}(G,\varphi)\neq\emptyset$, our algorithm recursively computes the vertex set $\mathsf{DP}_{w}[G,\varphi]$ that approximately achieves the maximum in Equation (1) or minimum in Equation (2). When it is clear from the context, we omit the subscript $w$ and simply write $\mathsf{opt}[G,\varphi]$ and $\mathsf{DP}[G,\varphi]$. We perform the dynamic programming algorithm using the formulas defined in Corollary 2.6 over the parse tree of a term obtained from Lemma 2.1 and Proposition 2.4. Specifically, at every stage of the algorithm, it is guaranteed that each $X\in\mathsf{sat}(G,\varphi)$ satisfies $X\subseteq V(G)\setminus\mathsf{src}(G)$. Furthermore, our algorithm is randomized. Thus, $\mathsf{DP}[G,\varphi]$ can be considered as a probability distribution over vertex sets in $\mathsf{sat}(G,\varphi)$.
@@ 3.3 Parallel Composition
$=\alpha\mathbb{E}\left[\mathsf{softmax}^{\varepsilon}_{i\in[p]}\mathsf{opt}_{i}\right]\geq(1-\varepsilon)\alpha\max_{i\in[p]}\mathsf{opt}_{i}=(1-\varepsilon)\alpha\mathsf{opt}[G\mathbin{//}H,\varphi].$
@@
where the last inequality is from $\varepsilon\leq 1$ and $\delta\leq\mathsf{opt}_{w}[G\mathbin{//}H,\varphi]$.
@@ 3.4 Forget
$=\alpha\mathbb{E}\left[\mathsf{softmax}^{\varepsilon}_{i\in[p]}\mathsf{opt}_{i}\right]\geq(1-\varepsilon)\alpha\max_{i\in[p]}\mathsf{opt}_{i}=(1-\varepsilon)\alpha\mathsf{opt}[\mathsf{fg}_{B}(G),\varphi].$
@@
$\leq 30\varepsilon^{-1}\log(2p\varepsilon^{-1})\delta+\beta+\delta\leq 31\varepsilon^{-1}\log(2p\varepsilon^{-1})\delta+\beta.$
@@ 3.5 Putting Together
Let $G=(V,E)$ be a graph with treewidth $k$. From Lemma 2.1 and Proposition 2.4, we can compute a HR-parse tree $t$ over $(k+1)$-graph denoting $G$ with height $O(\log k+\log n)\leq O(\log n)$. We perform the dynamic programming algorithm on this parse tree. We have the following.
Lemma 3.9.
@@
Proof 3.10.
The approximability bound is obtained by repeatedly applying Lemmas 3.1 and 3.5 and $(1-\varepsilon)^{h}\geq 1-h\varepsilon$. The Lipschitzness bound is obtained by repeatedly applying Lemmas 3.3 and 3.7.
@@
Proof 3.11 (Proof of Theorem 1.1).
The claim follows by substituting $\varepsilon$, $h$, and $k$ in Lemma 3.9 with $\frac{\varepsilon}{\Theta(\log n)}$, $O(\log n)$, and $23k+$ 2, respectively.
@@ 4.1 Independent Set
From the proof of Theorem 1.1, $f(k,\varphi)$ in Theorem 1.1 is bounded by $\log(2p_{\max})$. Particularly, if $p_{\max}$ is bounded by $2^{O(k)}$, the Lipschitz constant depends linearly on $k$. Here, we prove Theorem 1.3 for the maximum weight independent set problem by showing that $p_{\max}\leq 2^{O(k)}$. Similar arguments can be applied to several other problems, such as the minimum weight vertex cover problem and the minimum weight dominating set problem. For a $k$-graph $G$ and $S\subseteq\mathsf{src}(G)$, let
@@
Particularly, we have $p_{\max}\leq 2^{O(k)}$ and therefore Theorem 1.3 is proved.
@@ 5 Baker’s Technique
In this section, we provide a technique for Lipschitz continuous algorithms for optimization problems on planar graphs. Our technique is a modification of Baker’s technique [Baker94], which provides a PTAS for many optimization problems on planar graphs. For simplicity, in this section, we focus on the maximum weight independent set problem on planar graphs. Similar techniques can be applied to obtain Lipschitz continuous algorithms for other problems, such as the minimum weight vertex cover problem on planar graphs.
@@
Proof 5.4.
From Lemma 2.8, we have $\mathbb{E}\left[\mathsf{opt}_{\bm{j}^{*}}\right]\geq\left(1-\frac{1}{m}\right)\max_{j\in\{0,\dots,m-1\}}\mathsf{opt}_{j}$. For each $j\in\{0,\dots,m-1\}$, our framework outputs $\left(1-\frac{1}{m}\right)$-approximate solution for the maximum weight independent set problem on $G_{j}$. Therefore, the overall approximation ratio is $\left(1-\frac{1}{m}\right)^{2}\geq 1-\frac{2}{m}\geq 1-\varepsilon$.
Lemma 5.5.
@@
Proof 5.6.
Let $w\in\mathbb{R}_{\geq 0}^{V}$, $u\in V$, and $\delta>0$. We bound $\mathsf{EM}\left(\bm{X}_{w},\bm{X}_{w+\delta\bm{1}_{u}}\right)$, where $\bm{X}_{w}$ and $\bm{X}_{w+\delta\bm{1}_{u}}$ are the output of our algorithm on weight vector $w$ and $w+\delta\bm{1}_{u}$, respectively. We denote the variable $\bm{j}^{*}$ corresponding to the weight vector $w$ and $w+\delta\bm{1}_{u}$ by $\bm{j}^{*}$ and $\bm{j}^{*}_{w+\delta\bm{1}_{u}}$, respectively. From Lemma 2.9, we have
@@
$=O\left(\left(\varepsilon^{-1}+\log\log n\right)\varepsilon^{-1}\log^{2}n\right)\delta,$
where the first inequality is from Theorem 1.3.
@@ 6 Lower Bounds for Max Ones
In this section, we prove Theorem 1.5. To this end, we use a known sensitivity lower bound for the maximum cut problem. In the maximum cut problem, given a graph $G=(V,E)$, the goal is to find a bipartition of $V$ that maximizes the number of edges between the two parts. Then, the sensitivity of a randomized algorithm $\mathcal{A}$ for the maximum cut problem on a graph $G=(V,E)$ is defined to be
@@
Proof 6.2 (Proof of Theorem 1.5).
Let $\varepsilon,\delta>0$ be as in Lemma 6.1. Assume that there exists a $(1-\varepsilon)$-approximation algorithm $\mathcal{A}_{\rm MO}$ for the max ones problem with Lipschitz constant $s(n)$, where $n$ is the number of variables in the input 3CNF formula. Below, we design a low sensitivity algorithm $\mathcal{A}_{\rm MC}$ for the maximum cut problem using $\mathcal{A}_{\rm MO}$.
@@
This implies that $\mathcal{A}_{\rm MC}$ has sensitivity at most $O(s(n^{3}))$. By Lemma 6.1, we must have $s(n)=\Omega(n^{\delta/3})$.

### v2

@@ 2.3 Monadic Second-Order Logic
Then, we can design an algorithm for an (C)MSO maximization (minimization) problem over graphs of bounded treewidth. For simplicity, we assume that the given graph $G$ has no sources, that is, $\mathsf{src}(G)=\varnothing$. Let $t$ be a term denoting $G_{t}$ and $\varphi(X)$ be the (C)MSO-formula describing the constraint of the problem. We recursively compute the value $\mathsf{opt}(t^{\prime},\varphi^{\prime}\mathord{{\restriction}_{S}})=\max\{w(A):G\models\varphi^{\prime}\mathord{{\restriction}_{S}}(A)\}$ for any subterm $t^{\prime}$ of $t$ and MSO-formula $\varphi^{\prime}\mathord{{\restriction}_{S}}$ as follows, where $w\in\mathbb{R}_{\geq 0}^{V}$ is the given weight vector and $w(X)=\sum_{x\in X}w_{x}$ for any $X\subseteq V(G)$. If $t^{\prime}$ is of the form $\mathsf{fg}_{B}(t^{\prime\prime})$, then $\mathsf{opt}(\mathsf{fg}_{B}(t^{\prime\prime}),\varphi^{\prime}\mathord{{\restriction}_{S}})=\max_{S^{\prime}\subseteq B}\{\mathsf{opt}(t^{\prime\prime},\psi\mathord{{\restriction}_{S\cup S^{\prime}}})+w(S^{\prime})\}$, where, $\psi$ is the formula obtained from Corollary 14. If $t^{\prime}$ is of the form $t_{1}\mathbin{//}t_{2}$, then $\mathsf{opt}(t^{\prime},\varphi^{\prime}\mathord{{\restriction}_{S}})=\max_{i\in[p]}\{\mathsf{opt}(t_{1},\theta\mathord{{\restriction}_{S_{i}}})+\mathsf{opt}(t_{2},\psi\mathord{{\restriction}_{S^{\prime}_{i}}})\}$, where, $\{\theta_{i}\mathord{{\restriction}_{S_{i}}}({X}),\psi_{i}\mathord{{\restriction}_{S^{\prime}_{i}}}({X})\}_{i\in[p]}$ are the formulas obtained from Corollary 14. Then, $\mathsf{opt}(t,\varphi\mathord{{\restriction}_{\varnothing}})$ is the maximum value for the problem since $t$ has no sources. In the following, for simplicity, we may omit the $\mathord{{\restriction}_{S}}$ notation and consider only formulas $\varphi$ such that $\mathsf{sat}(G,\varphi)$ does not contain any sources.
In the above setting, the graphs are considered without any special vertex sets (called as colors) or special vertices (called as labels). However, for $k$-graphs with a constant number of colors and/or labels, an appropriate HR-algebra and logics can be defined, and the similar results for Corollary 14 hold (see e.g. [4, 18]).
@@ 2.4 Lipschitz Continuity
In our algorithm, we use the following procedures softmax and softmin. These procedures are derived from the exponential mechanism in the literature of differential privacy [8] and frequently appear in the literature on Lipschitz continuity [15, 17]. Here, we organize them into a more convenient form for our use. Let $p\in\mathbb{Z}_{\geq 1}$, $x_{1},\dots,x_{p}\in\mathbb{R}_{\geq 0}$, and $\epsilon\in(0,1]$. The softmax of $x_{1},\dots,x_{p}$ is taken as follows. If $\max_{i\in[p]}x_{i}=0$, we set $\mathsf{softargmax}^{\epsilon}_{i\in[p]}x_{i}$ to be an arbitrary index $i$ with $x_{i}=0$ and $\mathsf{softmax}^{\epsilon}_{i\in[p]}x_{i}=0$. Assume otherwise. First, we sample $\bm{c}$ uniformly from $\left[\frac{2\epsilon^{-1}\log(2p\epsilon^{-1})}{\max_{i\in[p]}x_{i}},\frac{4\epsilon^{-1}\log(2p\epsilon^{-1})}{\max_{i\in[p]}x_{i}}\right]$. Let $\bm{i}^{*}$ be a probability distribution over $[p]$ such that
@@
holds for all $i\in[p]$. Then, we define $\mathsf{softargmax}^{\epsilon}_{i\in[p]}x_{i}:=\bm{i}^{*}$ and $\mathsf{softmax}^{\epsilon}_{i\in[p]}x_{i}:=x_{\bm{i}^{*}}$. We have the following. The proofs of Lemmas 16 and 17 are given in Appendix B.
Lemma 16.
@@
holds for all $i\in[p]$. Then, we define $\mathsf{softargmin}^{\epsilon}_{i\in[p]}x_{i}:=\bm{i}^{*}$ and $\mathsf{softmin}^{\epsilon}_{i\in[p]}x_{i}:=x_{\bm{i}^{*}}$. We have the following. The proofs of Lemmas 18 and 19 are given in Appendix B.
Lemma 18.
@@ 3 Lipschitz Continuous Algorithms for $\textsf{MSO}_{2}$ Optimization Problems
In this section, we prove Theorems 1, 2 and 3. Since the proofs are similar to that for Theorem 1, we provide most of the proof of Theorem 2 in Appendix C. In Section 3.1, we define the notations and give a brief overview of the algorithm. Section 3.2 handles the base case, where the graph consists of a single vertex. Sections 3.3 and 3.4 analyze the impact of the parallel composition and the forget operations, respectively, on approximability and the Lipschitz constant. In Section 3.5, we put everything together to prove Theorem 1. Finally, in Section 4.1, we provide a more refined analysis for the special case of the maximum weight independent set problem to prove Theorem 3.
@@ 3.1 Definitions
for minimization problems. If $\mathsf{sat}(G,\varphi)=\emptyset$, we define $\mathsf{opt}_{w}[G,\varphi]=\mathtt{nil}$. For graphs $G$ and $\textsf{MSO}_{2}$ formulas $\varphi$ with $\mathsf{sat}(G,\varphi)\neq\emptyset$, our algorithm recursively computes the vertex set $\mathsf{DP}_{w}[G,\varphi]$ that approximately achieves the maximum in Equation (1) or minimum in Equation (2). When it is clear from the context, we omit the subscript $w$ and simply write $\mathsf{opt}[G,\varphi]$ and $\mathsf{DP}[G,\varphi]$. We perform the dynamic programming algorithm using the formulas defined in Corollary 14 over the parse tree of a term obtained from Lemma 8 and Proposition 11. Specifically, at every stage of the algorithm, it is guaranteed that each $X\in\mathsf{sat}(G,\varphi)$ satisfies $X\subseteq V(G)\setminus\mathsf{src}(G)$. Furthermore, our algorithm is randomized. Thus, $\mathsf{DP}[G,\varphi]$ can be considered as a probability distribution over vertex sets in $\mathsf{sat}(G,\varphi)$.
@@ 3.3 Parallel Composition
$=\alpha\mathbb{E}\left[\mathsf{softmax}^{\varepsilon}_{i\in[p]}\mathsf{opt}_{i}\right]\geq(1-\varepsilon)\alpha\max_{i\in[p]}\mathsf{opt}_{i}=(1-\varepsilon)\alpha\mathsf{opt}[G\mathbin{//}H,\varphi].\qed$
@@
where the last inequality is from $\varepsilon\leq 1$ and $\delta\leq\mathsf{opt}_{w}[G\mathbin{//}H,\varphi]$. ∎
@@ 3.4 Forget
$=\alpha\mathbb{E}\left[\mathsf{softmax}^{\varepsilon}_{i\in[p]}\mathsf{opt}_{i}\right]\geq(1-\varepsilon)\alpha\max_{i\in[p]}\mathsf{opt}_{i}=(1-\varepsilon)\alpha\mathsf{opt}[\mathsf{fg}_{B}(G),\varphi].\qed$
@@
$\leq 30\varepsilon^{-1}\log(2p\varepsilon^{-1})\delta+\beta+\delta\leq 31\varepsilon^{-1}\log(2p\varepsilon^{-1})\delta+\beta.\qed$
@@ 3.5 Putting Together
Let $G=(V,E)$ be a graph with treewidth $k$. From Lemma 8 and Proposition 11, we can compute a HR-parse tree $t$ over $(k+1)$-graph denoting $G$ with height $O(\log k+\log n)\leq O(\log n)$. We perform the dynamic programming algorithm on this parse tree. We have the following.
Lemma 24.
@@
Proof.
The approximability bound is obtained by repeatedly applying Lemmas 20 and 22 and $(1-\varepsilon)^{h}\geq 1-h\varepsilon$. The Lipschitzness bound is obtained by repeatedly applying Lemmas 21 and 23. ∎
@@
Proof of Theorem 1.
The claim follows by substituting $\varepsilon$, $h$, and $k$ in Lemma 24 with $\frac{\varepsilon}{\Theta(\log n)}$, $O(\log n)$, and $23k+$ 2, respectively. ∎
@@ 4.1 Independent Set
From the proof of Theorem 1, $f(k,\varphi)$ in Theorem 1 is bounded by $\log(2p_{\max})$. Particularly, if $p_{\max}$ is bounded by $2^{O(k)}$, the Lipschitz constant depends linearly on $k$. Here, we prove Theorem 3 for the maximum weight independent set problem by showing that $p_{\max}\leq 2^{O(k)}$. Similar arguments can be applied to several other problems, such as the minimum weight vertex cover problem and the minimum weight dominating set problem. For a $k$-graph $G$ and $S\subseteq\mathsf{src}(G)$, let
@@
Particularly, we have $p_{\max}\leq 2^{O(k)}$ and therefore Theorem 3 is proved.
@@ 5 Baker’s Technique
In this section, we provide a technique for Lipschitz continuous algorithms for optimization problems on planar graphs. Our technique is a modification of Baker’s technique [1], which provides a PTAS for many optimization problems on planar graphs. For simplicity, in this section, we focus on the maximum weight independent set problem on planar graphs. Similar techniques can be applied to obtain Lipschitz continuous algorithms for other problems, such as the minimum weight vertex cover problem on planar graphs.
@@
Proof.
From Lemma 16, we have $\mathbb{E}\left[\mathsf{opt}_{\bm{j}^{*}}\right]\geq\left(1-\frac{1}{m}\right)\max_{j\in\{0,\dots,m-1\}}\mathsf{opt}_{j}$. For each $j\in\{0,\dots,m-1\}$, our framework outputs $\left(1-\frac{1}{m}\right)$-approximate solution for the maximum weight independent set problem on $G_{j}$. Therefore, the overall approximation ratio is $\left(1-\frac{1}{m}\right)^{2}\geq 1-\frac{2}{m}\geq 1-\varepsilon$. ∎
Lemma 28.
@@
Proof.
Let $w\in\mathbb{R}_{\geq 0}^{V}$, $u\in V$, and $\delta>0$. We bound $\mathsf{EM}\left(\bm{X}_{w},\bm{X}_{w+\delta\bm{1}_{u}}\right)$, where $\bm{X}_{w}$ and $\bm{X}_{w+\delta\bm{1}_{u}}$ are the output of our algorithm on weight vector $w$ and $w+\delta\bm{1}_{u}$, respectively. We denote the variable $\bm{j}^{*}$ corresponding to the weight vector $w$ and $w+\delta\bm{1}_{u}$ by $\bm{j}^{*}$ and $\bm{j}^{*}_{w+\delta\bm{1}_{u}}$, respectively. From Lemma 17, we have
@@
$=O\left(\left(\varepsilon^{-1}+\log\log n\right)\varepsilon^{-1}\log^{2}n\right)\delta,\qed$
where the first inequality is from Theorem 3.
@@ 6 Lower Bounds for Max Ones
In this section, we prove Theorem 5. To this end, we use a known sensitivity lower bound for the maximum cut problem. In the maximum cut problem, given a graph $G=(V,E)$, the goal is to find a bipartition of $V$ that maximizes the number of edges between the two parts. Then, the sensitivity of a randomized algorithm $\mathcal{A}$ for the maximum cut problem on a graph $G=(V,E)$ is defined to be
@@
Proof of Theorem 5.
Let $\varepsilon,\delta>0$ be as in Lemma 29. Assume that there exists a $(1-\varepsilon)$-approximation algorithm $\mathcal{A}_{\rm MO}$ for the max ones problem with Lipschitz constant $s(n)$, where $n$ is the number of variables in the input 3CNF formula. Below, we design a low sensitivity algorithm $\mathcal{A}_{\rm MC}$ for the maximum cut problem using $\mathcal{A}_{\rm MO}$.
@@
This implies that $\mathcal{A}_{\rm MC}$ has sensitivity at most $O(s(n^{3}))$. By Lemma 29, we must have $s(n)=\Omega(n^{\delta/3})$. ∎